OpenStudy (anonymous):

find the nth partial sum of the arithmetic sequence: 1.50, 1.45, 1.40, 1.35 . . . n=20

OpenStudy (anonymous):

how do i figure it out???

OpenStudy (anonymous):

I send you the link it should be there

OpenStudy (anonymous):

what link

OpenStudy (anonymous):

It's just arithmetic, don't use "partial Sum" it confused the solvers: a20 = a1 + ( n -2) d = 1.5 + ( 19 ) .5 = 11

OpenStudy (anonymous):

Just plug the number in, the formula in your hand already!